# A Modular Approach for Multimodal Summarization of TV Shows

### v3

@@ 6 Conclusion
We addressed the task of summarizing TV shows from videos and dialogue transcripts. We proposed a modular approach where different specialized components perform separate sub-tasks. A scene-detection module splits the TV show into scenes, and a scene-reordering module places these scenes in an optimal order for summarization. A dialogue summarization module condenses the dialogue for each scene and a visual processing module produces a textual description of the video contents. Finally, a high-level summarization module fuses these into an output summary for the entire episode. We also introduced PREFS, a new metric for long-form summarization, based on splitting predicted and reference summaries into atomic facts. It captures both precision and recall, and correlates significantly with human judgments. In the future, we plan to test our method on even longer inputs, and explore settings where transcripts are not available.
@@
Our proposed PREFS metric requires multiple calls to GPT which incurs a financial cost. We estimate that all the results presented in this paper cost about \$1300. This is still many times cheaper than hiring human evaluators to extract and score facts manually, which we estimate would take 400 person-hours and cost about \$15,000.
There is still a significant gap, in terms of PREFS, between our summaries and the upper bound of comparing the gold summaries to each other. This shows that the task is challenging and requires further advances to reach human-level.
@@ Appendix A Example Summaries
Table 4 shows the gold summary for an episode from the SummScreen3D testset. Tables 5 and 68 show the summary of our model, and those of comparison models, respectively, for the same episode.
@@ Appendix B Caption Filtering and Inference of Names
Appendix C PREFS Metric
@@ E.1 BERTScore
Tables 10 and 11 report results on summary quality according to BERTScore Zhang et al. (2019). Tables 10 shows BERTScore for our model and comparison models, the analogue of Table 1 in the main paper, and Table 11 shows BERTScore for the ablation settings, the analogue of Table 2 in the main paper.
We find this metric does not distinguish well between the different settings, and scores all models very similarly. Even the “w/o transcript” setting, which qualitatively misses most of the important information in the episode, and scores poorly on ROUGE and PREFS, gets a high BERTScore. Moreover, even within each setting, the scores are very similar across different episodes, again in contrast to ROUGE, PREFS and qualitative evaluation (some episodes appear much easier to summarize than others). The average standard deviation across episodes, within each setting, is $2.2$, $1.39$, and $1.41$ for BERTScore-precision, BERTScore-recall and BERTScore-f1 respectively. In contrast, the same standard deviation for ROUGE-1, ROUGE-2, ROUGE-Lsum are $6.66$, $3.32$, $6.52$, while for factscore-precision, factscore-recall and PREFS, they are $14.92$, $15.14$ $14.94$, respectively.
@@ Appendix G Correlation Between Metrics
Figure 3 shows the pairwise correlation (Pearson) between all metrics that we report in Section 5. This is taken across all data points in all settings. All metrics are at least weakly correlated with each other, with the strongest correlations between the different varieties of ROUGE. This suggests that our PREFS metric captures information outside of what is captured by ROUGE.

### v4

@@ 6 Conclusion
We addressed the task of summarizing TV shows from videos and dialogue transcripts. We proposed a modular approach where different specialized components perform separate sub-tasks. A scene-detection module splits the TV show into scenes, and a scene-reordering module places these scenes in an optimal order for summarization. A dialogue summarization module condenses the dialogue for each scene and a visual processing module produces a textual description of the video contents. Finally, a high-level summarization module fuses these into an output summary for the entire episode. We also introduced PRisma, a new metric for long-form summarization, based on splitting predicted and reference summaries into atomic facts. It captures both precision and recall, and correlates significantly with human judgments. In the future, we plan to test our method on even longer inputs, and explore settings where transcripts are not available.
@@
Our proposed PRisma metric requires multiple calls to GPT which incurs a financial cost. We estimate that all the results presented in this paper cost about \$1300. This is still many times cheaper than hiring human evaluators to extract and score facts manually, which we estimate would take 400 person-hours and cost about \$15,000.
There is still a significant gap, in terms of PRisma, between our summaries and the upper bound of comparing the gold summaries to each other. This shows that the task is challenging and requires further advances to reach human-level.
Acknowledgments
We thank the anonymous reviewers for their feedback. We gratefully acknowledge the support of Amazon and the UK Engineering and Physical Sciences Research Council (grant EP/W002876/1), and the feedback and advice from Liz Norred, Zhu Liu, Vimal Bhat and Rohith Mysore Vijaya Kumar.
@@ Appendix A Example Summaries
Table 5 shows the gold summary for an episode from the SummScreen3D testset. Tables 6 and 79 show the summary of our model, and those of comparison models, respectively, for the same episode.
@@ Appendix B Caption Filtering and Inference of Names
Appendix C PRisma Metric
@@ E.1 BERTScore
Tables 11 and 12 report results on summary quality according to BERTScore Zhang et al. (2019). Tables 11 shows BERTScore for our model and comparison models, the analogue of Table 1 in the main paper, and Table 12 shows BERTScore for the ablation settings, the analogue of Table 2 in the main paper.
We find this metric does not distinguish well between the different settings, and scores all models very similarly. Even the “w/o transcript” setting, which qualitatively misses most of the important information in the episode, and scores poorly on ROUGE and PRisma, gets a high BERTScore. Moreover, even within each setting, the scores are very similar across different episodes, again in contrast to ROUGE, PRisma and qualitative evaluation (some episodes appear much easier to summarize than others). The average standard deviation across episodes, within each setting, is $2.2$, $1.39$, and $1.41$ for BERTScore-precision, BERTScore-recall and BERTScore-f1 respectively. In contrast, the same standard deviation for ROUGE-1, ROUGE-2, ROUGE-Lsum are $6.66$, $3.32$, $6.52$, while for factscore-precision, factscore-recall and PRisma, they are $14.92$, $15.14$ $14.94$, respectively.
@@ Appendix G Correlation Between Metrics
Figure 3 shows the pairwise correlation (Pearson) between all metrics that we report in Section 5. This is taken across all data points in all settings. All metrics are at least weakly correlated with each other, with the strongest correlations between the different varieties of ROUGE. This suggests that our PRisma metric captures information outside of what is captured by ROUGE.
@@
Appendix H Human Evaluation Setup
We request that each participant read the following guidelines before ranking the summaries.
When comparing summaries, you should take into account the following three points, in order of importance. (1) Firstly, most importantly, does it give plenty of detail about the events in the episode, rather than repeating the same information multiple times? (2) Does it report events and characters that really are in the episode, rather than irrelevant information, e.g. about when and where the episode was aired or reviews it received? (3) Is the text well-written, containing full properly-formed sentences devoid of spelling, grammar, and punctuation errors?
The following is an example of a good summary:
Clint goes to Robert’s home and has two guys assault him for what he did to Jessica. But he tells Robert he still gets to keep the job that Clint has enabled him to keep at LU and James can keep his grant. Meanwhile, Jessica and Brody are happily planning their future together with their new baby but agree to get the paternity test so that they know. In the place where Eli has kept her, Dani manages to untie herself from the chair and rushes to the other room in an attempt to save Starr and Hope, assuming they are on the other side. But when she gets through, she is shocked and elated to see her mom. They cry and rejoice but are horrified to find out that Eli, whom they’ve both seen as a trusted family member and upstanding citizen, is a serial killer and psycho. She admits she does not remember much but is going to get better and be with her daughter and husband. Dani unties her mom and they are ready to go home while Eli is gone. Eli goes to find unconscious Greg in his hospital room and is ready to inject a lethal dose and kill him knowing what Greg was planning, until Shaun catches him and he runs. Greg is still unconscious however, and Eli escapes. And he comes back in time to prevent Tea and Dani from escaping. Meanwhile, Hannah holds Starr and baby Hope captive and is determined to keep Starr away from Cole for as long as it takes for him to forget all about Starr and fall in love with Hannah. But when she returns to Marty’s, James notices that she is carrying his bullet on the chain in her purse. He finds that very odd since it was last in Starr’s possession. He also finds it very odd when Nate informs him and Robert that Eli called and let him, Todd and Blair all talk to Dani but made no mention of Starr and nobody knows where Starr and Hope are. He senses they are in danger and is determined to save them.
The following is an example of a bad summary, which repeats the same point again and again instead of covering all the information in the episode:
Ford is beaten up by two men that Clint hires, but Clint does not fire him from his job at NU, even though Ford expects that he will be fired. Clint is angry because Ford took advantage of his daughter, Jessica, while she was vulnerable. Jessica and Brody are worried about who the father of their baby is, so they are doing a paternity test. They are doing the paternity test so they can find out who the father is.They are worried about who the father is, but they are going to do a paternity test so they can find out who the father is.They are worried about who the father is, but they are going to do a paternity test so they can find out who the father is.The paternity test will tell them who the father is, and they are worried about it because then they will find out who the father is. They are doing the paternity test to find this out, and find out who the father is with the paternity test. They are very worried about it, but are doing a paternity test to find out for sure. This will tell them who the father is.
The following is an example of a bad summary that is poorly written, discusses irrelevant material and contains garbled sentences:
Clint pays two large men to come with him to Ford’s apartment and beat him up for what Ford@; Ford did, the apartment downstairs, Ford did to his daughter. Ford thinks he will be fired from NU but Clint says he is not. Ford: I guess I’m gonna be out of a job, huh? At The The.. the hospital, Brody meets Jessica to do a paternity test. .. Brody and.. The hospital. The test is ready to go. The episode airs at 9 p.m. ET/PT on Thursday, December 14. The series finale airs on December 15 at 8 p.M. ET on ABC. Jessica: whatever happens, we’ll raise this baby together,. Brody: I’m glad you think so honey. Jessica: let’s just do the test to be sure. Dani: Mom? I never thought I’d find you here, I thought you were dead. Tea: I’m so happy that you’re here. Now let’s get out of here before Eli comes back. The season finale airs at 10 p. M.E. on December 16 at 8:30 p.S. ET.M ET on E.C. ET at the Palace Hotel in New York City, New York, and on December 18 at 9 a.C.. The finale will air at 9 P.M.. The season. The finale airs. The premie.. The premieThe finale he series.. The season finale. The.. season finale..The. season finale aired.. The. finale airs.. September 14 at 9:00 p. m. ET, September 14, 2013,
The ‘good summary‘ example is a gold summary from the train set. The ‘bad summary‘ examples are composed by us.
We also ask each participant a question about the contents the episode they are assigned, to check that they have watched and understood the full episode. The list of episodes, and corresponding questions and answers are shown in Table 13.
Episode Question Answer
Port Charles, 05-06-03 what sketch does Kevin show to Elizabeth? a sketch of her.
The Bold and the Beautiful, 05-22-06 what does Stephanie promise to never do again if Brooke comes back to the company? disrespect her
Guiding Light, 02-28-05 what flavour ice cream did the grandson, Zach, get all over his face? chocolate fudge swirl
One Life to Live, 01-06-12 What are the names of Allison’s two daughters that Vicki faked the paternity test for? Natalie and Jessica
The Bold and the Beautiful, 04-27-06 what is Stephanie trying to convince Bridget not to do? take her grandson to Italy
Guiding Light, 09-06-06 who is Tammy in love with? Johnathan
As the World Turn, 01-06-05 who does Henry say he’s in love with? Katie
Guiding Light, 01-24-05 who does Harley think Beth might have done? Killed Phillip
The Bold and the Beautiful, 01-02-15 who is Rick convincing to sign the papers? Eric, his dad
Table 13: The questions, and corresponding answers, for each episode that we ask particpants in order to check they fully watched and understood the video. Episode names are in the format <show-name> <date-aired (MM-DD-YYYY)>.